### v2

# TUNI: A Textual Unimodal Detector for Identity Inference in CLIP Models

Abstract
The widespread usage of large-scale multimodal models like CLIP has heightened concerns about the leakage of PII. Existing methods for identity inference in CLIP models require querying the model with full PII, including textual descriptions of the person and corresponding images (e.g., the name and the face photo of the person). However, applying images may risk exposing personal information to target models, as the image might not have been previously encountered by the target model. Additionally, previous MIAs train shadow models to mimic the behaviors of the target model, which incurs high computational costs, especially for large CLIP models. To address these challenges, we propose a textual unimodal detector (TUNI) in CLIP models, a novel technique for identity inference that: 1) only utilizes text data to query the target model; and 2) eliminates the need for training shadow models. Extensive experiments of TUNI across various CLIP model architectures and datasets demonstrate its superior performance over baselines, albeit with only text data.
TUNI: A Textual Unimodal Detector for Identity Inference in CLIP Models
Songze Li1,∗,†, Ruoxi Cheng1,thanks: Contributed equally to this work. 1Southeast University, Nanjing China. 2Nanyang Technological University, Singapore. Corresponding authors: songzeli@seu.edu.cn. , Xiaojun Jia2
1 Introduction
Recent years have witnessed a rapid development of large-scale multimodal models, such as Contrastive Language–Image Pre-training (CLIP) (Radford et al., 2021). These models synthesize information across different modalities, particularly text and images, facilitating applications from automated image generation to sophisticated visual question answering systems. Despite their potential, these models pose significant privacy risks (Inan et al., 2021; Carlini et al., 2021; Leino and Fredrikson, 2020; Rigaki and Garcia, 2023; Helbling et al., 2023; Rahman et al., 2024; Rahman, 2023) as the vast datasets used for training often contain personally identifiable information (PII) (Schwartz and Solove, 2011; Abadi et al., 2016; Bonawitz et al., 2017), raising concerns (Xi et al., 2024) about PII leakage and misuse (Hu et al., 2023; Yin et al., 2021). Therefore, it is extremely important to develop tools to detect potential PII leakage from CLIP models. Specially, as the first step, we would like to address the identity inference problem, i.e., to determine if the PII of a particular person was used in training of a target CLIP model.
Figure 1: Current methods query LLMs with both text and image, while our goal is to conduct identity inference with only textual data.
Traditional methods, like Membership Inference Attacks (MIAs) (Shokri et al., 2017), have focused on determining whether a specific data sample was used for model training. When applied to CLIP models, these approaches typically involve querying the model with both texts and images of the target individual (Ko et al., 2023), and exposing images of a person the CLIP model may have not seen in the training set brings new privacy leakage risk (He et al., 2022). Hence, it is desirable to have a detection mechanism for ID inference that does not query the CLIP model with real images of the person (see an example in Figure 1). Furthermore, traditional MIAs often rely on constructing shadow models that mimic the behaviors of the target model to obtain training data to construct attack models  (Hu et al., 2022a), which demands extensive computational resources and is less feasible in environments with limited computational capabilities (Mattern et al., 2023; Hisamoto et al., 2020; Jagielski et al., 2024). Alternative methods for shadow models in MIAs, such as those based on cosine similarity (Ko et al., 2023) and self-influence functions (Cohen and Giryes, 2024), exhibit either lower accuracy or still necessitate substantial computational resources (Oh et al., 2023).
Figure 2: Features of textual descriptions extracted from the optimized images guided by a CLIP model with ResNet50x4 architecture, trained on a dataset where each person has 75 images. The cosine similarity between the embeddings of optimized image and the tested text, and the distance between the embeddings of the optimized images, can clearly distinguish between the samples within and outside the training dataset of the target CLIP model.
To address these limitations, we propose a textual unimodel detector (TUNI) for identity inference in CLIP models, which queries the target model with only text information during inference. Specifically, we first propose a feature extractor, which maps a textual description to a feature vector through image optimization guided by the CLIP model; then, we randomly generate a large amount of textual gibberish, which we know do not match any textual descriptions in the training dataset. As shown in Figure 2, we make the key observation that the feature distributions of textual gibberish and member samples in the training set are well distinguishable.
Leveraging this property, we use the feature vectors of the generated textual gibberish to train multiple anomaly detectors to form an anomaly detection voting system. At test time, TUNI simply feeds the feature vector of the test text to the voting system, and determines that if the corresponding PII is included in the training set (abnormal) or not (normal). The training of the anomaly detector in TUNI costs only several hours with four NVIDIA GeForce RTX 3090 GPUs, avoiding training shadow models with the size of the CLIP model in traditional MIAs, which can cost over 18 days even with hundreds of advanced GPUs (Gu et al., 2022; Ko et al., 2023; Hu et al., 2022b).
Our contributions are summarized as follows:
  • We propose a textual unimodal detector, dubbed TUNI, which is the first method to conduct identity inference in CLIP models with unimodal data, preventing risky exposure of images to the target model;
  • We find that the feature distributions of texts that are in and out of the target CLIP model are well separated, and propose to adopt randomly generated text to train anomaly detectors for ID inference, avoiding the need for computationally intensive shadow models in traditional MIAs.
  • Extensive experiments conducted across six kinds of CLIP models have indicated that the proposed TUNI achieves better performance than current methods for identity inference, even when using only textual data.
2 Related Work
2.1 Privacy Leakage in CLIP Models
CLIP model exemplifies modern multimodal innovation by integrating an image encoder and a text encoder into its architecture (Radford et al., 2021). These encoders transform inputs into a shared embedding space, enabling effective measurement of semantic similarity (Ramesh et al., 2022). Despite the significant advances and expansive applicability of CLIP models, the vast and diverse datasets utilized for training such models could potentially include sensitive information, raising concerns about privacy leakage (Hu et al., 2022b). Various inference attacks, including model stealing  (Dziedzic et al., 2022; Liu et al., 2022; Wu et al., 2022), knowledge stealing (Liang et al., 2022), data stealing (He and Zhang, 2021), and membership inference attacks (Liu et al., 2021; Ko et al., 2023), have been developed for CLIP, exposing potential vulnerability in privacy leakage. These privacy concerns underscore the necessity for developing robust defense mechanisms to safeguard sensitive information in CLIP models (Golatkar et al., 2022; Jia et al., 2023; Huang et al., 2023).
2.2 Personally Identifiable Information and Leakage Issues
Personally Identifiable Information (PII) is defined as any data that can either independently or when combined with other information, identify an individual. Training Large Language Models (LLMs) often utilizes publicly accessible datasets, which may inadvertently contain PII. This elevates the risk of data breaches that could compromise individual privacy and entail severe legal and reputational consequences for the deploying entities (Lukas et al., 2023; Abadi et al., 2016; Bonawitz et al., 2017; Rahman et al., 2020; Shamshad et al., 2023). Various attacks have been developed to reveal PII from LLMs. A method is proposed in (Panda et al., 2024) to steal private information from LLMs via crafting specific queries to GPT-4 that can reveal sensitive data by appending a secret suffix to the generated text; Zhang et al. introduced the ETHICIST method for targeted training data extraction, through loss smoothed soft prompting and calibrated confidence estimation, significantly improving extraction performance on public benchmarks (Zhang et al., 2023); Carlini et al. also studied training data extraction from LLMs, emphasizing the predictive capability of attacks given a prefix (Carlini et al., 2021); ProPILE, proposed in (Kim et al., 2024), probes privacy leakage in LLMs, by assessing the leakage risk of PII included in the publicly available Pile dataset; Inan et al. investigated the risks associated with membership inference attacks using a Reddit dataset, further emphasizing the persistent threat of PII leakage in various data environments (Inan et al., 2021).
2.3 Current Identity Inference Methods and Their Limitations
Identity inference, critical in privacy-preserving data analysis, has garnered significant attention across domains, such as genomic data (Erlich et al., 2018), location-based spatial queries (Kalnis et al., 2007), person re-identification scenarios (Karaman and Bagdanov, 2012), computer-mediated communication (Motahari et al., 2009)and face recognition (Zhou and Lam, 2018; Prince et al., 2011; Sanderson and Lovell, 2009). Membership Inference Attacks (MIAs), which determine if specific data points were in a model’s training dataset, can be used to perform identity inference. Traditional MIAs often require constructing shadow models to mimic the target model’s behavior, posing computational efficiency challenges for large models (Truex et al., 2019; Ye et al., 2022; Meeus et al., 2023; Xue et al., 2023).
While identity inference has been mainly performed on unimodal models, it is recently extended to CLIP models. Identity Detection Inference Attack (IDIA) (Hintersdorf et al., 2022) does not need shadow models; it involves providing real photos of the tested individual and 1000 prompt templates including the real name to choose from. The attacker generates multiple queries by substituting the <NAME> placeholder and analyzes the model’s responses to calculate an attack score based on correct predictions. If the correct name is predicted for a threshold number of templates, the individual is inferred to be in the training data. Cosine Similarity Attacks (CSA) (Ko et al., 2023) uses cosine similarity (CS) between image and text features to infer membership, as CLIP is trained to maximize CS for training samples. Based on CSA, Weak Supervision Attack (WSA) uses a new weak supervision MIA framework with unilateral non-member information for enhancement. Both IDIA and WSA avoid the high costs associated with shadow models, but require querying the target model with real images the model may have never seen, raising new privacy concerns.
3 Methodology
3.1 Problem Setup and Threat Model
Consider a CLIP model $M$ trained on a dataset $D_{\text{train}}$. Each sample $s_{i}=(t_{i},x_{i})$ in $D_{\text{train}}$ records the personally identifiable information (PII) of an individual person, and consists of a textual description $t_{i}$ (e.g., name of the person) and a corresponding image $x_{i}$ (e.g., face photo of the person). For distinct indices $i\neq j$, it is possible that $t_{i}=t_{j}$ and $x_{i}\neq x_{j}$, indicating that multiple non-identical images of the same person may exist.
A detector would like to probe potential leakage of a person’s PII through the target CLIP model $M$, via conducting an identity inference task against $M$, to determine if any PII samples of this person were included in the training set $D_{\text{train}}$.
Detector’s Goal. For a person with textual description $t$, a detector would like to determine whether there exits a PII sample $(t_{i},x_{i})\in D_{\text{train}}$, such that $t_{i}=t$.
Note that rather than detecting for a particular text-image pair $(t,x)$, our goal is to detect existence of any (one or more) pair with a textual description of $t$. This is because that multiple images of the same person can be used for training, and any one of these images may lead to potential PII leakage.
Detector’s Knowledge and Capability. The detector can query $M$ and observe the output, including extracted image and text embeddings as well as their matching score, but does not know the model architecture of $M$, the parameter values, or the training algorithms. For the target textual description $t$, depending on the application scenarios, the detector may or may not have actual images corresponding to $t$. Nevertheless, in the case where the detector knows corresponding images, due to privacy concerns, it cannot include them in the queries to $M$. The detector cannot modify $M$ or access its internal state.
3.2 TUNI: Textual Unimodal Detector for ID Inference
We design a textual unimodal detector for ID inference (TUNI), to determine whether the PII of a person is in the training set of the target CLIP model $M$, with the restriction that only the textual description of the person can be exposed to $M$. Firstly, for a textual description $t$, we develop a feature extractor to map $t$ to a feature vector, through image optimization guided by the CLIP model. Then, we make the key observation that textual gibberish like “D2;l-NOXRT”—random combinations of numbers and symbols clearly do not match any textual descriptions in the training set, and hence the detector can generate large amount of textual gibberish that are known out of $D_{\text{train}}$. Using feature vectors extracted from these textual gibberish, the detector can train multiple anomaly detectors to form an anomaly detection voting system. Finally, during the inference phase, the features of the target textual description are fed into the system, and the inference result is determined through voting. Additionally, when the actual images of the textual description is available to the detector, they can be leverage to perform clustering on the feature vectors of the test samples to further enhance detection performance. An overview of the proposed TUNI framework is shown in Figure 3.
Figure 3: Overview of TUNI.
Feature Extraction through CLIP-guided Image Optimization. The feature extraction for a textual description $t$ involves iterative optimization of an image $x$, to maximize the correlation between the embeddings of $t$ and $x$ out of the target CLIP model. The extraction process, described in Algorithm 1, iterates for $n$ epochs; and within each epoch, an image is optimized for $m$ iterations, to maximize the cosine similarity between its embedding of the CLIP model and that of the target textual description. The average optimized cosine similarity $S$ and standard deviation of the optimized image embeddings $D$ are extracted as the features of $t$ from model $M$.
Algorithm 1 CLIP-guided Feature Extraction
Input: Target CLIP model $M$, textual description $t$
Output: Mean optimized cosine similarity $S$, standard deviation of optimized image embeddings $D$
1:$n\leftarrow$ number of epochs
2:$m\leftarrow$ number of optimization iterations per epoch
3:$\mathcal{S}\leftarrow\emptyset$, $\mathcal{V}\leftarrow\emptyset$
4:$v_{t}\leftarrow M(t)$triangleright Obtain text embedding from $M$
5:for $i=1$ to $n$ do
6:     $x_{0}\leftarrow\textup{Rand}()$triangleright Randomly generate an initial image
7:     for $j=0$ to $m-1$ do
8:         $v_{x_{j}}\leftarrow M(x_{j})$triangleright Obtain image embedding from $M$
9:         $x_{j+1}\leftarrow\arg\max_{x_{j}}\frac{v_{t}\cdot v_{x_{j}}}{\|v_{t}\|\,\|v_{x_{j}}\|}$triangleright Update image to maximize cosine similarity
10:     end for
11:     $S_{i}\leftarrow\frac{v_{t}\cdot v_{x_{m}}}{\|v_{t}\|\,\|v_{x_{m}}\|}$triangleright Optimized similarity for epoch $i$
12:     $\mathcal{S}\leftarrow\mathcal{S}\cup\{S_{i}\}$, $\mathcal{V}\leftarrow\mathcal{V}\cup\{v_{x_{m}}\}$
13:end for
14:$S\leftarrow\frac{1}{n}\sum_{S_{i}\in\mathcal{S}}S_{i}$
15:$\bar{v}\leftarrow\frac{1}{n}\sum_{v\in\mathcal{V}}v$
16:$D\leftarrow\sqrt{\frac{1}{n}\sum_{v\in\mathcal{V}}\|v-\bar{v}\|^{2}}$
17:return $S$, $D$
Generation of Textual Gibberish. TUNI starts the detection process with generating a set of $\ell$ gibberish strings $\mathcal{G}=\{g_{1},g_{2},\ldots,g_{\ell}\}$, which are random combinations of digits and symbols with certain length. As these gibberish texts are randomly generated at the inference time, with overwhelming probability that they did not appear in the training set. Applying the proposed feature extraction algorithm on ${\cal G}$, we obtain $\ell$ feature vectors =${\cal F}=\{f_{1},f_{2},\ldots,f_{\ell}\}$ of the gibberish texts.
Training Anomaly Detectors. Motivated by the observations in Figure 2 that the feature vectors of the texts that are in and out of the training set of $M$ are well separated, we propose to train an anomaly detector using ${\cal F}$, such that texts out of $D_{\text{train}}$ are considered “normal”, and the problem of ID inference on textual description $t$ is converted to anomaly detection on the feature vector of $t$. More specifically, $t$ is detected to be in $D_{\text{train}}$, if its feature vector is detected “abnormal” by the trained anomaly detector. Specifically in TUNI, we train several anomaly detection models on ${\cal F}$, such as Isolation Forest, LocalOutlierFactor (Cheng et al., 2019) and AutoEncoder (Chandola et al., 2009). These models constitute an anomaly detection voting system that will be used for ID inference on the test textual descriptions.
Textual ID Inference through Voting. For each textual description $t$ in the test set, TUNI first extracts its feature vector $f$ using Algorithm 1, and then feeds $f$ to each of the obtained anomaly detectors to cast a vote on whether $t$ is an anomaly. When the total number of votes exceeds a predefined detetion threshold $N$, $t$ is determined as an anomaly, i.e., PII with textual description $t$ is used to train the CLIP model $M$; otherwise, $t$ is considered normal and no PII with $t$ is leaked through training of $M$.
Table 1: Performance comparison with baseline methods across different CLIP models. $\Delta$ indicates the improvement of TUNI.
Architecture Number of photos per Method Precision $\Delta$ Recall $\Delta$ Accuracy $\Delta$
person in training set
ResNet-50 1 WSA 0.6653 ± 0.0032 0.1979 0.2925 ± 0.0045 0.6896 0.6675 ± 0.0037 0.2497
IDIA 0.6922 ± 0.0023 0.1712 0.4032 ± 0.0027 0.5789 0.6836 ± 0.0034 0.2336
TUNI 0.8634 ± 0.0031 - 0.9821 ± 0.0042 - 0.9172 ± 0.0028 -
75 WSA 0.6625 ± 0.0018 0.2017 0.2867 ± 0.0061 0.6968 0.6710 ± 0.0043 0.2322
IDIA 0.6901 ± 0.0024 0.1741 0.3998 ± 0.0049 0.5837 0.6907 ± 0.0075 0.2125
TUNI 0.8642 ± 0.0057 - 0.9835 ± 0.0019 - 0.9032 ± 0.0033 -
ResNet-50x4 1 WSA 0.6712 ± 0.0029 0.1901 0.2912 ± 0.0048 0.6835 0.6808 ± 0.0031 0.2547
IDIA 0.6625 ± 0.0036 0.1963 0.3980 ± 0.0031 0.5267 0.6957 ± 0.0029 0.2398
TUNI 0.8613 ± 0.0033 - 0.9747 ± 0.0013 - 0.9355 ± 0.0038 -
75 WSA 0.6724 ± 0.0022 0.1988 0.2935 ± 0.0054 0.6981 0.6685 ± 0.0047 0.2777
IDIA 0.7085 ± 0.0021 0.1627 0.3904 ± 0.0018 0.6012 0.7167 ± 0.0035 0.2295
TUNI 0.8712 ± 0.0043 - 0.9916 ± 0.0037 - 0.9462 ± 0.0029 -
ViT-B/32 1 WSA 0.6323 ± 0.0064 0.0268 0.2964 ± 0.0052 0.3421 0.6812 ± 0.0045 0.0025
IDIA 0.6783 ± 0.0047 0.0308 0.3746 ± 0.0033 0.2639 0.6772 ± 0.0041 0.0065
TUNI 0.7091 ± 0.0056 - 0.6385 ± 0.0062 - 0.6837 ± 0.0044 -
75 WSA 0.7045 ± 0.0075 0.0137 0.2806 ± 0.0048 0.3566 0.6895 ± 0.0052 0.0052
IDIA 0.6890 ± 0.0051 0.0292 0.3811 ± 0.0063 0.2561 0.6927 ± 0.0045 0.0020
TUNI 0.7182 ± 0.0068 - 0.6372 ± 0.0046 - 0.6947 ± 0.0078 -
Enhancement with Real Images. At inference time, if real images of the test texts are available at the detector (e.g., photos of a person), they can be used to extract an additional feature measuring the average distance between the embeddings of real images and those of optimized images using the CLIP model, using which the feature vectors of the test texts can be clustered into two partitions with one in $D_{\text{train}}$ and another one out of $D_{\text{train}}$. This adds an additional vote for each test text to the above described anomaly detection voting system, potentially facilitating the detection accuracy.
Specifically, for each test text $t$, the detector is equipped with a set of $c$ real images $\{x_{\mathrm{real}}^{1},x_{\mathrm{real}}^{2},\ldots,x_{\mathrm{real}}^{c}\}$. Similar to the feature extraction process in Algorithm 1, over $k$ epochs with independent initializations, $k$ optimized images $\{x_{\mathrm{opt}}^{1},x_{\mathrm{opt}}^{2},\ldots,x_{\mathrm{opt}}^{k}\}$ for $t$ are obtained under the guidance of the CLIP model. Then, we apply a pretrained feature extraction model $F$ (e.g., DeepFace (Taigman et al., 2014) for face images) to the real and optimized images to obtain real embeddings $\{v_{\mathrm{real}}^{1},v_{\mathrm{real}}^{2},\ldots,v_{\mathrm{real}}^{c}\}$ and optimized embeddings $\{v_{\mathrm{opt}}^{1},v_{\mathrm{opt}}^{2},\ldots,v_{\mathrm{opt}}^{k}\}$. Finally, we compute average pair-wise $\ell_{2}$ distance between real and optimized embeddings, denoted by $R$, over $c\cdot k$ pairs, and use $R$ as an additional feature of the text $t$.
For a batch of $B$ test texts $(t_{1},t_{2},\ldots,t_{B})$, we start with extracting their features $((S_{1},D_{1},R_{1}),(S_{2},D_{2},R_{2}),\ldots,(S_{B},D_{B},R_{B}))$. Feeding the first two features $S_{i}$ and $D_{i}$ into the trained anomaly detection system, each text $t_{i}$ obtains an anomaly score as the number of anomaly detectors who believe that it is abnormal. Additional, the $K$-means algorithm with $K=2$ is performed on the feature vectors $\{(S_{i},D_{i},R_{i})\}_{i=1}^{B}$ to partition them into a “normal” cluster and an “abnormal” cluster, adding another vote on the anomaly score of each test instance. Then, the ID inference of each text is performed by comparing its total number of received votes and a detection threshold $N^{\prime}$.
4 Evaluations
We evaluate the performance of TUNI, for the task of ID inference from the name of a person, with the corresponding image being the face photo of the person.
4.1 Setup
Our experiments leverage datasets and target CLIP models from (Hintersdorf et al., 2022).
Figure 4: Samples from the dataset for training CLIP models.
Dataset Construction. The datasets for training and ID inference are constructed from three datasets: LAION-5B (Schuhmann et al., 2022), Conceptual Captions 3M (CC3M) (Changpinyo et al., 2021), and FaceScrub (Kemelmacher-Shlizerman et al., 2016). Specifically, 200 celebrities—100 for training and 100 for validation, with their face photos accompanied by labels containing their names are selected from the FaceScrub dataset; then these data samples are augmented by additional photos of the selected celebrities found in LAION-5B, such that each person has multiple photos; finally these augmented data points are mixed with the CC3M dataset to form the training set of the CLIP model. By doing this, we have the ground truth on which people are in the training set and which are not. In our experiments, we construct two datasets, one with a single photo for each person, and another with 75 photos for each person. Samples of this dataset are shown in Figure 4 and a more detailed description is given in appendix.
Models. Our analysis involves ID inference from six pre-trained target CLIP models, categorized into ResNet-50, ResNet-50x4, and ViT-B/32 architectures. The ResNet-50 and ResNet-50x4 models are based on the ResNet architecture (He et al., 2016; Theckedath and Sedamkar, 2020); and ViT-B/32 models employ the Vision Transformer architecture (Chen et al., 2021). DeepFace (Serengil and Ozpinar, 2020) is used for facial feature extraction for enhancement with real images.
Evaluation Metrics. TUNI’s effectiveness is assessed using Precision, Recall, and Accuracy metrics, measuring anomaly prediction accuracy, correct anomaly identification, and overall prediction correctness, respectively.
Baselines. Current ID inference detection methods for CLIP models typically require detector to query target model with corresponding real images. Most MIAs involve training shadow models and related methods like shadow encoders  (Liu et al., 2021), which can be particularly costly for large-scale multimodal models. We empirically compare the performance of TUNI with the following SOTA inference methods, which both avoid using shadow models, but still require submitting both text and image to the target CLIP model for inference.
  • Identity Inference Attack (IDIA) (Hintersdorf et al., 2022) detects with a list of 1000 names to choose from and 30 real photos for a tested person. In IDIA, the attacker (detector) selects candidate names as prompt templates, and predicts names for each image and prompt. Once the correct name is predicted, it’s inferred that the target individual is in training dataset. We compare IDIA using 3 photos for each test sample with TUNI using only text.
  • Weakly Supervised Attack (WSA) (Ko et al., 2023) uses cosine similarity between image and text features to infer membership, and adds a weak supervision MIA framework based on non-member data generated after the release of the target model.
All experiments are performed using four NVIDIA GeForce RTX 3090 GPUs. Each experiment is repeated for 10 times, and the average values and the standard deviations are reported.
4.2 Results
On training anomaly detectors, we randomly generated $\ell=50$ textual gibberish (some of them are shown in Table 3).
The image optimization was performed for $n=100$ epochs; and in each epoch, $m=1000$ Gradient Descent (GD) iterations with a learning rate of $0.02$. Four anomaly detection models, i.e., LocalOutlierFactor (Cheng et al., 2019), IsolationForest (Liu et al., 2008), OneClassSVM (Li et al., 2003; Khan and Madden, 2014), and AutoEncoder (Chen et al., 2018) were trained, and $N=3$ was chosen as the detection threshold.
As shown in Table 1, TUNI, even with only text information, consistently outperforms WSA and IDIA in all metrics by a large margin, across all model architectures and datasets, demonstrating its superior performance.
We also evaluate the effect of providing the TUNI detector with an real photo of the inferred person. In this case, the embedding distances between the real and optimized images of the test samples are used to perform a $2$-means clustering, adding another vote to the inference result. We accordingly raise the detection threshold $N^{\prime}$ to $4$. As illustrated in Table 2, the given photo helps to improve the performance of TUNI across all tested CLIP models. While recalls in some ResNet models experience minor declines attributed to the raised threshold, all remain above 94%. Conversely, the ViT-B models exhibit an almost 11% increase in recall. A lower detection threshold aids recall enhancement but may concurrently lead to declines in other metrics.
Table 2: Detection performance with a given photo during inference. $\Delta$ indicates performance improvement.
Architecture Number of photos per TUNI Precision $\Delta$ Recall $\Delta$ Accuracy $\Delta$
person in training set
ResNet-50 1 Text only 0.8634 ± 0.0031 0.1019 0.9821 ± 0.0042 -0.0396 0.9172 ± 0.0028 0.0303
With 1 photo 0.9653 ± 0.0032 - 0.9425 ± 0.0057 - 0.9475 ± 0.0041 -
75 Text only 0.8642 ± 0.0057 0.1183 0.9835 ± 0.0019 -0.0188 0.9032 ± 0.0033 0.0538
With 1 photo 0.9825 ± 0.0031 - 0.9467 ± 0.0024 - 0.9570 ± 0.0038 -
ResNet-50x4 1 Text only 0.8613 ± 0.0033 0.1290 0.9747 ± 0.0013 -0.0183 0.9355 ± 0.0038 0.0317
With 1 photo 0.9923 ± 0.0011 - 0.9564 ± 0.0044 - 0.9672 ± 0.0028 -
75 Text only 0.8712 ± 0.0043 0.0912 0.9916 ± 0.0037 0.0019 0.9462 ± 0.0029 0.0323
With 1 photo 0.9624 ± 0.0042 - 0.9935 ± 0.0029 - 0.9785 ± 0.0037 -
ViT-B/32 1 Text only 0.7091 ± 0.0056 0.1432 0.6385 ± 0.0062 0.1084 0.6837 ± 0.0044 0.0975
With 1 photo 0.8523 ± 0.0038 - 0.7469 ± 0.0078 - 0.7812 ± 0.0031 -
75 Text only 0.7182 ± 0.0068 0.1353 0.6372 ± 0.0046 0.1086 0.6947 ± 0.0078 0.1148
With 1 photo 0.8535 ± 0.0042 - 0.7458 ± 0.0039 - 0.8095 ± 0.0063 -
4.3 Ablation Study
We further explore the impacts of different system parameters on the detection accuracy.
Figure 5: Detection accuracy for different numbers of optimization iterations per epoch.
Figure 6: Detection accuracy for different numbers of epochs.
Figure 7: Detection accuracy with name only.
Figure 8: Detection accuracy with a face photo.
Figure 9: Detection accuracy for different numbers of gibberish.
Figure 10: Detection accuracy for different number of real photos.
Optimization parameters. Figure 6 and 6 show that during feature extraction, optimizing for $n=100$ epochs, each with $m=1,000$ iterations, offers the optimal performance. Additional epochs and optimization iterations, while incurring additional computational cost, do not significantly improve the detection accuracy.
Detection threshold. Figure 8 and 8 show that the system attains higher accuracy, when it adopts a threshold of three votes for considering an input as an anomaly with text only, and four votes with an added detection model using an additional given photo. Setting a high threshold may result in failing to detect an anomaly, while setting a low one may lead to identifying a normal one as anomaly.
Number of textual gibberish. As shown in Figure 10, for different target models, the detection accuracies initially improve as the number of gibberish texts increases, and converge after using more than $50$ gibberish strings.
Number of real photos. As shown in Figure 10, integrating real photos can enhance the detection accuracy; however, the improvements of using more than 1 photo are rather marginal.
Table 3: Samples of randomly generated gibberish.
+̱7IKXb2Y FR!pnI<5xS euiT_;yw/
je|%5(s=G\_ ?Ŵ<E{Dvmz hqf- =j<q5
#lEZ0yrZ5ig ’2_:6[jiOa X*|<tFx|4/
Fa<Z*Oike[ \93W4̃>x5u ?=&QplxC-c
5 Defense and Covert Gibberish Generation
In real-world scenarios, target models being detected may deploy defense mechanisms to recognize anomalous inputs like gibberish and provide misleading outputs, causing TUNI to misjudge inclusion of PII.
To generate more covert gibberish data, we can create strings resembling normal text, with a few characters replaced by syllables from another language. For instance, the detector can craft query texts, by randomly combining English names with syllables from Arabic medical terminology. One way to do this is to start by prompting LLMs like GPT-3.5-turbo to create lists of common initial and final syllables in English words. These syllable lists are then extracted and refined to ensure diversity and eliminate duplicates. Next, the refined syllable combinations are randomly paired to create pseudo-English names, such as “Karinix”, “Zylogene”, “Glycogenyx”, and “Renotyl”. It’s crucial to verify the novelty of these names by checking against a database of real names to avoid collision. Then by prompting the LLM to generate strings using the refined syllable combinations, covert gibberish strings resembling real names are produced (some examples are given in Table 4).
Table 4: Covert gibberish that seem to be real names.
Karinix Zylogene Glycogenyx
Zylotrax Vexilith Dynatrix
Exodynix Novylith Glycosyne
Xenolynx Rynexis Delphylith
6 Conclusion
In this paper, we propose TUNI, the first method to conduct identity inference without exposing acutal images to target CLIP models. TUNI turns inference problem into anomaly detection, through randomly generating textual gibberish that are known to be out of training set, and exploting them to train anomaly detectors. Furthermore, the incorporation of real images is shown to enhance detection performance. Through evaluations across various CLIP model architectures and datasets, we demonstrate the consistent superiority of TUNI over baselines.
7 Limitations
Due to constraints resources, we conducted experiments using the name of the individual as textual descriptions. This approach may not fully encapsulate the complexities and nuances of real-world PII leakage including addresses, phone numbers, and other sensitive information.
8 Ethics and Social Impact
The development of TUNI highlights crucial ethical considerations in identity inference using multimodal models like CLIP. By enabling identity inference with only textual data, TUNI reduces the risks associated with exposing PII through images. This approach not only helps protect individual privacy but also minimizes the potential for misuse in harmful applications. As such technologies evolve, it is essential for researchers to adhere to ethical guidelines and promote transparency, ensuring that advancements in AI prioritize user privacy and foster responsible usage in society.
9 Potential Risks
TUNI aims to bolster privacy by aiding in identity inference and safeguarding personal identifiable information within AI systems. While mindful of the risk of misuse, TUNI should adhere to data regulations and be employed only with explicit consent from involved data subjects, promoting privacy and security in AI practices.
Appendix A Dataset Description
We utilized the datasets from previous work (Hintersdorf et al., 2022).
LAION-400M (Schuhmann et al., 2021), comprising 400 million image-text pairs, primarily employed for pre-training the CLIP model, offering a wide array of visual content and textual descriptions to facilitate the model’s learning of relationships between images and text, including direct associations between specific individuals and images. In the experiment, this dataset is used to analyze the frequency of individuals appearing within it to identify individuals with lower frequencies of appearance, thereby avoiding the use of those individuals that appear very frequently to prevent skewing the experimental results. A threshold is set to only use individuals with fewer than 300 appearances for the experiments to ensure that the experimental results would not be dominated by individuals with very high occurrence frequencies, thus ensuring the accuracy and reliability of the experimental outcomes.
LAION-5B (Schuhmann et al., 2022), containing over 5.8 billion pairs and LAION-400M is its subset. In the experiment, LAION-5B is used to expand the CC3M dataset, enriching and increasing the sample size and diversity of the dataset. LAION-5B is used to find similar pairs to those in the FaceScrub dataset for each of the 530 celebrities. After confirming the presence of these celebrities’ names in the captions of the found images, these image-text pairs were added to the CC3M dataset for training the target CLIP models.
Conceptual Captions 3M (CC3M) (Changpinyo et al., 2021), consisting of 2.8 million image-text pairs, anonymizes image captions by replacing named entities (e.g., celebrity names) with their hypernyms (e.g., "actor"). This dataset was also employed for pre-training the CLIP model. However, in this experiment, researchers analyzed the dataset using facial recognition technology to determine if specific celebrity images were present, and selectively added image-text pairs for model training adversarial attacks. As the named entities in CC3M dataset are anonymized in image captions, i.e., specific celebrity names replaced with their hypernyms like "actor," after confirming the presence or absence of specific celebrity images in the CC3M dataset, controlled additions of image-text pairs were made to the CC3M dataset.
FaceScrub (Kemelmacher-Shlizerman et al., 2016), containing images of 530 celebrities, was used to ascertain whether the identities one intends to infer are part of the training data. Celebrities were chosen due to the wide availability of their images in the public domain, minimizing privacy concerns associated with using their images.
To accurately calculate evaluation metrics, it was necessary to analyze which individuals were already part of the dataset and which were not. For the LAION-5B dataset, names of the 530 celebrities from the FaceScrub dataset were searched within all captions, and corresponding image-text pairs were saved, which were then added to the CC3M dataset. This was done to train the CLIP model and evaluate the effectiveness of IDIA under controlled conditions. In the experiments with the CC3M dataset, a total of 200 individuals were used, with 100 added to the dataset for model training and the remaining 100 held out for model validation. The selection of data in this process was balanced in terms of gender, with an equal distribution of male and female individuals to enhance the persuasiveness of the results. We construct two datasets for training the CLIP models of three architectures relatively, one with a single photo for each person, and another with 75 photos for each person. Samples of the datasets are shown in Figure 4.